the frequency of a paticle performing a shm is 12hz its amplitude is 4cm its initial displacement is 2cm towards positive extreme position its equation of displacement is
 

Grade:11

1 Answers

Arun
6 years ago
Dear Aditya
 
Equation of displacement is given by
X = A sin(wt + @)
Where x = final displacement
 
Given that x= 2 (initial displacement)
f = 12 Hz 
w = 2 * pi * f  = 24 \pi rad /sec
Amplitude A = 4 cm
We have to find angle @,
x = A sin@
2 = 4 sin@
Sin@ = ½
@ = \pi/6
 
Hencev equation of displacement is
X = 4 sin(24\pi t + \pi/6)
 
Regards
Arun (askIITians forum expert)